OpenStudy (anonymous):

integral cos^(3)xsin^(4)x dx. Need help breaking this down to integrate.

OpenStudy (ash2326):

Problem will become easy, if we could change its form to something like \[\int\sin^ n x\times \cos x\] Now if we put sin x = t, cos x dx=dt It'll become \[\int t^n dt\]

OpenStudy (ash2326):

Let's try that \[\cos^3 x \sin ^ 4x\] \[(1-\sin^2 x )\cos x \times \sin^4 x\] \[\sin^4 x \times \cos x -\sin^6 x \times \cos x\] Do you get this?

OpenStudy (anonymous):

yes i do

OpenStudy (ash2326):

Now, just substitute sin x = t and then integrate. Can you do that?

OpenStudy (anonymous):

so you would be looking something like:\[t ^{4}*cosx-t ^{6}*cosx\] is that what it would look like before i Integrate?

OpenStudy (ash2326):

\[\int (t^4 -t^6 )dt\]

OpenStudy (ash2326):

Have you studied substitution method?

OpenStudy (anonymous):

no, it is an online class and they are jumping around.

OpenStudy (ash2326):

Ok, I'll explain

OpenStudy (ash2326):

Suppose we have the simple integral \[\int \sin x\times\cos x\ dx\] There are several ways of doing this, I'll use the substitution method Let \[\sin x = t \]Now differentiate it, with respect to x \[\cos x =\frac {dt}{dx}\] or \[\cos x\ dx= dt\] Our integral is \[\int \sin x\times\cos x\ dx\] substituting sin x =t and cos x dx =dt We get \[\int t\ dt\] Integration is easy \[\frac {t^2} 2 +c\] Put the value of t back \[\frac{\sin^2x}{2}+C\]

OpenStudy (anonymous):

wow that is simple

OpenStudy (ash2326):

In you problem, after using substitution you'll get \[\int (t^4-t^6 ) dt\] Remember we have substituted sin x =t Not sin^4 x =t or sin^6 x =t If we do that, there are consequences :P

OpenStudy (anonymous):

so i get the following.\[((5sinx)/5)-((7sinx)/7)+c\] does that appear correct?

OpenStudy (ash2326):

\[\frac{\sin^5 x }{5}-\frac{\sin^7 x}{7}+C\] Use latex to type equations, I've used the following code \frac{\sin^5 x }{5}-\frac{\sin^7 x}{7}+C

OpenStudy (anonymous):

ok so the only issue i have is the answer I am supposed to get is \[\frac{ 1 }{ 70 }\sin ^{5}(x)(5\cos(2x)+9)+c\]

OpenStudy (ash2326):

We can get that as well. Pull out sin^ 5 x you'll get a sin^2 x term, substitute \[\sin^ 2 x =\frac{1-\cos 2 x}{2}\] Then simplify

OpenStudy (anonymous):

thank you, I think I got it.

OpenStudy (ash2326):

Welcome :)